indias mangaiyan was sent to the Mars by launching it into a tranfer orbit EOMaround the sum it leaves the earth at E andmeets Mars at M IF THE SEMI MAJOR AXIS OF EARTHS ORBIT IS ae=150000000km that of Mars orbit am=228000000km taken  kelers laws give the estimate of time for Mangalyanto reach Mars from earth to be closed to

Grade:12

1 Answers

Arun
6 years ago
Dear Prasad
 
r = (1.5 + 2.28) / 2 = 1.89
 
Tm / Te = (1.89 / 1.5)3/2
 
tm = Tm / 2
 = (365/2) * 1.41
 = 257.3 days = 260 days
 
Regards
Arun (askIITians forum expert)